OpenStudy (anonymous):

the length of a rectangle is 2 inches longer than its width. if the perimeter is 36 inches, find the length and width.

OpenStudy (shrutipande9):

perimeter is 2(l+w) and l=w+2

OpenStudy (anonymous):

let w=width, l=length, P=perimetre P=2w+2l l=2+w P=2w+2(2+w) P=4w+4 36=4w+4 32=4w w=8 l=2+8 l=10

OpenStudy (shrutipande9):

i hope u can solve further...:)